A 0.600-kg ball is dropped from rest at a point 4.50 m above the floor. The ball rebounds straight upward to a height of 3.40 m. Taking the negative direction to be downward, what is the impulse of the net force applied to the ball during the collision with the floor?
Answer #1

velocity with which it strikes the ground

v^2 = 2 gh

v^2 = 2*9.8* 4.5

v = 9.39 m/s

velocoty with which it rebounds

u^2 = 2 g h' = 2*9.8* 3.4

u = 8.613 m/s

Impulse

I = m ( u - v)

I = 0.6* ( 9.39 + 8.613)

I = 10.532 kg m/s

=======

do comment in case any doubt, will reply for sure. 5